OpenStudy (anonymous):

suppose f and g are functions for which the lim f(x)=25 as x goes to -2 and lim g(x) =10 use limit properties to find the following limits. a. lim g(x)/f(x) as x goes to -2 b. lim square root f(x)g(x) as x goes to -2 c. lim 4f(x)+2/1-g(x)+f(x) as x goes to -2 Someone please help ! Thanks

OpenStudy (anonymous):

suppose f and g are functions for which the lim f(x)=25 as x goes to -2 and lim g(x) =10 use limit properties to find the following limits a.\[\lim_{x \rightarrow -2}g(x) \div f(x)\] b. \[\lim_{x \rightarrow -2}\sqrt{f(x)g(x)}\] c. \[\lim_{x \rightarrow -2}4f(x)+2 \div 1-g(x)+f(x)\]

OpenStudy (anonymous):

a. 10/-2=-5 b. sqrt -20 = (2i)sqrt5 c. 4(-2) +2/1-10 -2=-4/11

OpenStudy (anonymous):

how did you work be out

OpenStudy (anonymous):

sqrt -20 =(sqrt -1)(sqrt20)=(sqrt -1)(sqrt4*5)=i*2sqrt5

OpenStudy (anonymous):

O still don't get it.

OpenStudy (anonymous):

sqrt -1= i

OpenStudy (anonymous):

ok